A woman is one more year than 4 times the age of her daughter. In 4 years, she will be 3 times as old as her daughter will be then. What are their present ages?
Answer by rossiv53(27)

Let x = daughter's age and (4x+1) = the mother's age. In 4 years their ages will be (x+4) = daughter's age and (4x+1)+4 = the mother's age. Now in 4 years the mother will be 3 times as old. This means the equation is 3 times the daughter's age in 4 years = to the mother's age in 4 years.
3(x+4) = (4x+1)+4)
3x+12 = 4x+5
3x+12-12 = 4x+5-12
3x = 4x-7
3x-4x = -7
-x = -7
x = 7
So, daughter's age now is 7 and since mother's age is 4x+1, her age is 29.
Good luck!